Units:
NA= 6.02×1023 atoms/mol Walter White who lives in Simeranya (a utopic city in Anarres) prepares blue crystals, named as myth, which adopt a side-cent
Answer #1

Question 1.

The molecular formula of the myth can be written as follows.

A-atoms occupy all the 12 edge-centers.

The contribution by each edge-center = 1/4

Therefore, no. of atoms 'A' = 12*1/4 = 3

B-atom occupies the body-center.

The contribution by body-center = 1

Therefore, no. of atoms 'B' = 1*1 = 1

Therefore, the formula of the myth = A3B

Question 2.

The atomic packing factor (APF) of the unit cell = (Z*4/3 \pir3)/a3

Where Z = no. of atoms per unit cell = 3 + 1 = 4

'r' is the radius of the unit cell

'a' is the edge length of the unit cell

'a3' is the total volume of the unit cell

Question 3.

The density of the unit cell (d) = ZM/NA.a3

Here, M = molar mass of A3B = 3*58 + 60 = 234 g/mol

NA is Avagadro's number = 6.02*1023 atoms/mol

'a' is the edge-length of the unit cell

Question 4.

The coordination number of B = 12